OpenStudy (anonymous):

hat time is 2:45 ET in pacific time ?

OpenStudy (anonymous):

what

OpenStudy (amistre64):

california is 3 hours behind florida

OpenStudy (amistre64):

360/24 = 180/12 = 90/6 = 15 15 degrees is 1 hour so cali is about 45degrees from florida :)